OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sine of seven pi divided by eight. This is my answer below. Can someone please check it to see if I have the correct answer step by step. sin (7π/8)=√1-cos(7π/4/2) sin (7π/4)=√2/√2 sin Θ/2=±√1-cos (Θ)/2 =±√1-√2/2/2 +√2-√2/4=+√2-√2/2=2/√2-2

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[\large \sin \left(\dfrac{7\pi}{8}\right) = \sqrt{\dfrac{1-\cos(7\pi/4)}{2}}\]

ganeshie8 (ganeshie8):

cos(7pi/4) = cos(pi/4) = 1/sqrt(2)

ganeshie8 (ganeshie8):

\[\large \begin{align}\sin \left(\dfrac{7\pi}{8}\right) &= \sqrt{\dfrac{1-\cos(7\pi/4)}{2}}\\~\\&=\sqrt{\dfrac{1-1/\sqrt{2}}{2}}\\~\\&=\sqrt{\dfrac{\sqrt{2}-1}{2\sqrt{2}}}\\~\\&= \sqrt{\dfrac{2-\sqrt{2}}{4}}\\~\\&= \dfrac{\sqrt{2-\sqrt{2}}}{2}\end{align}\]

ganeshie8 (ganeshie8):

btw, your answer is also correct

OpenStudy (anonymous):

Is that a different way to do it or is my process wrong?

ganeshie8 (ganeshie8):

yeah i see few mistakes in your work in last two lines now, but its hard to make sense of what you wrote without proper latex :O

OpenStudy (anonymous):

Okay and thank you :)